The amplitude of a damped osscilator is become half on one minute.The amplitude after 3 minute will be 1/X times the original where X is

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Pallavi

Equation of under damped vibration is given as

y=Ye-ξωnt sin(wdt +Φ )

where is a amplitude at perticular time

at t=1 sec


Y/2 =Ye-ξωn

or 1/2 =e-ξωnt


now at t= 3 sec

Y'=Ye-ξωn3

   =Y(1/2)3

    =Y/8

 so X=8 answer


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin